OpenStudy (anonymous):

What is the value of 4x + 7 + 2y, if x= 6 and 7 = 4?

OpenStudy (anonymous):

Plug in 6 for x and 4 for y in the equation.

OpenStudy (anonymous):

It would be 39 right

OpenStudy (anonymous):

Right :)

OpenStudy (anonymous):

Thnks;) a hace another question

OpenStudy (anonymous):

What is it?

OpenStudy (anonymous):

What is the value of the expression 1/3 t + 2/3r if t =9 and r =6?

OpenStudy (anonymous):

Plug in 9 for t and 6 for r in the equation. 1/3(9) + 2/3(6).

OpenStudy (anonymous):

I know i dont get tht one :( my ópticos r 5, 7 or 9 but i cant find the anwered :(

OpenStudy (anonymous):

Okay, let's break it down. What is 9/3?

OpenStudy (anonymous):

Instead of trying to find (1/3) * 9, you can covert 9 into a fraction and multiply. 9 as a fraction is simply 9/1. \[\frac{ 1 }{ 3 } * \frac{ 9 }{ 1 } = ?\]

OpenStudy (anonymous):

Dont get it sorry:(

OpenStudy (anonymous):

What part is confusing you?

OpenStudy (anonymous):

Ok the result is suppose to be A. 5 B. 7 or C. 9 and I'm not getting the result

OpenStudy (anonymous):

Okay, let's go back tot he beginning and just look at the equation. \[\frac{ 1 }{ 3 }t + \frac{ 2 }{ 3 } r\] Plug in the values for t and r, which are given to you. t = 9, r = 6. \[\frac{ 1 }{ 3 }(9) + \frac{ 2 }{ 3 } (6)\] There are a couple ways to figure this out, but the easiest would be to convert the 9 and the 6 into fractions. You know what 9 as a fraction is \[\frac{ 9 }{ 1 }\] because 9/1 = 9. Then we can simply multiply. \[\frac{ 1 }{ 3 } * \frac{ 9 }{ 1 }\] Multiplying fractions is very easy. You just multiply the top numbers, then you multiply the bottom numbers. So let's multiply the top numbers first. The top numbers are 1 and 9. 1*9 = 9. Now let's multiply the bottom numbers. The bottom numbers are 3 and 1. 3*1 = 3. Now let's go back and write that as a fraction again.\[\frac{ 9 }{ 3 }\]. What is 9 divided by 3?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

But that's not in the answered tht I can choose

OpenStudy (anonymous):

That's only the first part of the equation. So 1/3(9) = 3. Now let's look at the second part of the equation. 2/3(6). You have to convert 6 into a fraction, which is simply 6/1 because 6/1 = 6. \[\frac{ 2 }{ 3 } * \frac{ 6 }{ 1 }\] Multiply the top numbers, which are 2 and 6. 2 * 6 = 12. Now multiply the bottom numbers, which are 3 and 1. 3 * 1 = 3. Now let's write that as a fraction. \[\frac{ 12 }{ 3 }\] What is 12/3?

OpenStudy (anonymous):

Is 4 but is not in there is either 5 , 7 or 9 :(

OpenStudy (anonymous):

So 1/3(9) + 2/3(6). We know that 1/3(9) = 3 and 2/3(6) = 4. So what's 3 + 4?